OpenStudy (anonymous):

x-xnaught=vnaught + (1/2)at^2 Naught is like x0 (starting position) how was the above equation derived? Please explain in simplest terms. math proof not really necessary. but logic will suffice.

OpenStudy (anonymous):

a is acceleration v is velocity

OpenStudy (anonymous):

actually vnaught (t) srry!

ganeshie8 (ganeshie8):

\(\large x - x_0 = v_0t + \frac{1}{2}at^2\)

OpenStudy (anonymous):

right, ok

ganeshie8 (ganeshie8):

u hav some familiarity wid calculus ? dx/dt, d^x/dt^2 etc...

OpenStudy (anonymous):

a bit like derivatives, chain rules, product rule,

ganeshie8 (ganeshie8):

that will do :) lets analyze the given equation

ganeshie8 (ganeshie8):

it looks like a quadratic in t ?

ganeshie8 (ganeshie8):

\(\large x - x_0 = v_0t + \frac{1}{2}at^2\) \(\large x = x_0 + v_0t + \frac{1}{2}at^2\)

ganeshie8 (ganeshie8):

notice that \(x\) is a quadratic function in \(t\),

ganeshie8 (ganeshie8):

u must be knowing that \(x\) is position of object, and \(t\) is time

ganeshie8 (ganeshie8):

lets take a simple quadratic equation and try to derive above

OpenStudy (anonymous):

sorry my internet disconnected for a while!

ganeshie8 (ganeshie8):

its okay :)

ganeshie8 (ganeshie8):

we want formula for position of one-dimensional motion wid constant acceleration. so we better take a quadratc :- \(\large x = at^2 + bt + c\)

ganeshie8 (ganeshie8):

thats the position function. how do we get VELOCITY function from above ?

ganeshie8 (ganeshie8):

\(\large v = \frac{dx}{dt}\) right ? so simply take derivative of position function , \(x\)

ganeshie8 (ganeshie8):

\(\large v = \frac{dx}{dt} = \frac{d}{dt} (at^2+ bt + c)\)

ganeshie8 (ganeshie8):

wat do u get ?

OpenStudy (anonymous):

2at+b right?

ganeshie8 (ganeshie8):

exactly, thats the velocity of object

OpenStudy (anonymous):

y-yo=1/2(a)(t-to)^2+vo(t-to)

ganeshie8 (ganeshie8):

\(\large x = at^2+ bt + c\) --------(1) \(\large v = \frac{dx}{dt} = \frac{d}{dt} (at^2+ bt + c) = 2at + b\) ----(2)

ganeshie8 (ganeshie8):

find acceleration now

ganeshie8 (ganeshie8):

\(\large a = \frac{dv}{dt} = \frac{d}{dt} (2at+b) = ?\)

OpenStudy (anonymous):

that would be the derivative of 2at+b which is 2a??

ganeshie8 (ganeshie8):

yup !

ganeshie8 (ganeshie8):

\(\large x = at^2+ bt + c\) --------(1) \(\large v = \frac{dx}{dt} = \frac{d}{dt} (at^2+ bt + c) = 2at + b\) ----(2) \(\large acc = \frac{dv}{dt} = \frac{d}{dt} (2at+b) = 2a\) ----(3)

ganeshie8 (ganeshie8):

from equation (3), can u solve the coefficient a ?

OpenStudy (anonymous):

before u use the equation u need to think about position and time that it start

ganeshie8 (ganeshie8):

from (3) \(\large acc = 2a\) =>? \(\large a = \frac{acc}{2}\) ----------------(4)

ganeshie8 (ganeshie8):

can u find the remaining coefficients, b and c

ganeshie8 (ganeshie8):

once we have values for , a, b and c, we can simplyy substitute this in (1) and get the equation for one-dimensional motion wid constant acc

ganeshie8 (ganeshie8):

thats a good point :) lets get to that after finishing playing wid quadratic @sovannayen

OpenStudy (anonymous):

b=v-At and c .. this is sort of complicated..

ganeshie8 (ganeshie8):

yes, put \(t=0\), then \(b\) becomes velocity at t=0 eh ?

OpenStudy (anonymous):

why is t 0?

ganeshie8 (ganeshie8):

put t= 0 in equation (2), \(\large v_0 = b\)

ganeshie8 (ganeshie8):

because, we want to derive final equaiton in terms of \(v_0\), which is the velocity at t=0

OpenStudy (anonymous):

oh I see..

ganeshie8 (ganeshie8):

im sure that doesnt convince u much, let me think of explaining u better

ganeshie8 (ganeshie8):

or, u can help me lol why t=0 above ?

OpenStudy (anonymous):

no I think I actually do understand. v naught equals the velocity at starting time. except, there should be a velocity at starting time shouldn't there?

OpenStudy (anonymous):

*shouldn't

ganeshie8 (ganeshie8):

there will be starting velocity for any object,

ganeshie8 (ganeshie8):

starting velocity can be 0, when its freeling falling from a height,

OpenStudy (anonymous):

but I don't think there is a velocity, in other words, there is a 0 velocity at starting time right?

ganeshie8 (ganeshie8):

but when u throw a stone, it surely will have a starting velocity other than 0

ganeshie8 (ganeshie8):